# The Potential and Challenges of Evaluating Attitudes, Opinions, and Values in Large Language Models

### v2

@@ 4.4 Evaluation
By drawing upon well-known measures of inter-annotator agreement and similarity measures, alignment metrics have been employed to measure the alignment of human and LLM responses. These measures include Cohen’s Kappa Argyle et al. (2023); Hwang et al. (2023), 1-Wasserstein distance (WD) Santurkar et al. (2023); Hwang et al. (2023); Sanders et al. (2023), Kullback–Leibler (KL) divergence Dominguez-Olmedo et al. (2023); Sun et al. (2024), the Euclidean distance between the model’s responses and the standard scores of humans Wang et al. (2023b), Jensen-Shannon Distance for model and country alignment Durmus et al. (2023), as well as correlation and statistical analysis Kalinin (2023); Sun et al. (2024); Jiang et al. (2024). Moreover, metrics have been applied to measure the alignment between variables, such as regression models for measuring the correlations between single features of different personas Bisbee et al. (2023) and between different nations Benkler et al. (2023).
@@
Several studies conduct robustness evaluations on LLM outputs to validate the opinion consistency under different prompt perturbations or persona steerings, typically by measuring the percentage of samples that reach the same answer as a consistency score (e.g. Feng et al., 2023; Hartmann et al., 2023; Röttger et al., 2024; Shu et al., 2024), or by measuring the entropy of answers (e.g. Dominguez-Olmedo et al., 2023; Tjuatja et al., 2024; Wang et al., 2024b). A few more specific metrics for robustness evaluation have been applied or put forward. Santurkar et al. (2023) introduce a consistency metric by measuring whether the LLM consistently aligns with the same group across different topics. Bonagiri et al. (2024) propose a metric named Semantic Graph Entropy (SaGE) to assess LLMs’ consistency level while answering moral questions with the same meaning but paraphrased in different formats. Wright et al. (2024) employ the average total variation distance (TVD) between the model responses across different demographic categories to validate the answer distances.
5 Opportunities and Challenges in Evaluating AOV in LLMs
Drawing from findings summarized in §3 and §4 from recent advances, we now focus on the methodological and practical perspectives regarding opportunities and challenges of evaluating AOV in LLMs (WHERE?). This section addresses several key issues starting with the need to understand the models themselves (step 1), followed by the necessity for human-AI alignment (step 2), and finally, the implications from and for downstream applications in social sciences (step 3).
@@ 5.1 Understanding the Model
The essential discussion on the impact of evaluating AOV in LLMs should start with the models themselves – the agents creating output. Our understanding of these models is limited (much like our understanding of ourselves as humans) Hassija et al. (2024). As studying how people respond to questions and express opinions helps us understand human behavior, examining how models do the same can enhance our knowledge of these models.
Evaluating AOV Helps Understand Model Behavior.
By effectively evaluating AOV in LLMs, we could potentially better explain their behavior in those subjective contexts, which could reveal why models produce certain opinions and values, helping us to better interpret their outputs. Apart from the textual output, tracking model internal behavior is also of interest, for example, to examine whether there exist skill neurons Wang et al. (2022); Voita et al. (2023). Investigating the internal working mechanisms of models enhances their interpretability, helping to make their operations more transparent and understandable. Currently, there is a lack of work linking AOV evaluations to model interpretability. Addressing this gap would significantly contribute to the understanding and reliability of LLM outputs, especially in subjective contexts.
Evaluating AOV Helps Understand Model Biases.
@@
One challenge in evaluating the output of LLMs is that the methods used can themselves be brittle. For example, in multiple-choice survey question settings, several studies rely on the first token logits (probabilities) of model output to map the options with the highest logits (as discussed in the Logits-Based Output section in §4.3). However, Wang et al. (2024a, b) observe that the first token logits do not always match the textual outputs and sometimes the mismatch rate can be over 50% in Llama2-7B Touvron et al. (2023) and Gemma-7B Team et al. (2024). A few works have also highlighted models being sensitive to option ordering Binz and Schulz (2023); Pezeshkpour and Hruschka (2024); Zheng et al. (2024); Shu et al. (2024); Wei et al. (2024), or generating inconsistent outputs to semantically equivalent situations Jang and Lukasiewicz (2023); Bonagiri et al. (2024); Wright et al. (2024), making evaluation unstable. Therefore, any evaluation for AOV in LLMs should be accompanied by extensive robustness tests (see the Robustness Metrics section in §4.4). Wang et al. (2024a, b) propose to look at the text by training classifiers on the annotated LLM outputs, which typically requires a lot of human effort and may not be generalizable. Developing context-aware evaluation metrics to capture human-like nuances in LLM outputs is an ongoing research focus for model interpretability.
@@ 5.2 Human-AI Alignment
After understanding the model, aligning LLMs with human AOV and ensuring that they perform safely and effectively is the next crucial phase.
@@
Alignment methods, such as Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022), focus on the problem of aligning LLMs to human values, which requires transferring the human values into alignment target for training and evaluating the models Klingefjord et al. (2024). However, current evaluations have often been coarse, highlighting the need for more fine-grained benchmarks to assess alignment effectively Lee et al. (2024b). One fundamental challenge RLHF faces is the problem of misspecification Casper et al. (2023). The diversity of human values cannot be easily represented by a single reward function. Current alignment evaluation benchmarks and reward model training rely on individual preference but lack consideration of the nature of diversity in human opinion. A more fine-grained evaluation of AOV with respect to social choice Conitzer et al. (2024) or social awareness Yang et al. (2024) will help us better understand the alignment process and design a better socially-aware alignment algorithms Conitzer et al. (2024).
@@ 5.3 Implications from and for Social Science Applications
Considering the potential and challenges from the model perspective, we will now explore the feasibility of deploying AOV in LLMs in downstream social science applications. LLMs, with their ability to process vast amounts of text data, could provide valuable insights into human values and behaviors. Again, caution must be exercised to address inherent biases and alignment issues that may arise.
@@
Currently, we have no means of aligning LLMs to accurately represent the diversity of human opinions necessary for reliable public opinion polling and similar tasks. The existing literature highlights numerous challenges, particularly in replicating non-US values Benkler et al. (2023); Arora et al. (2023); Simmons (2023); Rao et al. (2023). While some argue that LLM surveys might provide insights into hard-to-reach populations, the risk remains significant that these groups are difficult to model accurately by LLMs von der Heyde et al. (2023); Namikoshi et al. (2024).
Human AOV Help Evaluate AOV in LLMs.
While there is a great gap between Human AOV and those in LLMs, human-centered applications can enhance our understanding and validation of AOV in LLMs. In survey methodology, responding to a survey question involves several cognitive steps, mainly including comprehension, retrieval, judgment, and reporting Tourangeau et al. (2000); Groves et al. (2004); Tourangeau (2018). Figure 2 illustrates a basic model of the human survey response process. Despite fundamental differences, the behavioral study of machines can benefit from that of animals Rahwan et al. (2019), as well as of humans Greasley and Owen (2016). Therefore, by integrating these human-centered cognitive processes into the examination of how LLMs respond to survey questions, we are able to gain valuable insights into the models and then modify the models to better align with human processes. Still, while concepts from human AOV are certainly helpful in studying LLMs, we should also keep in mind at all times that they are after all not humans and should caution against the anthropomorphism we discussed in the previous section.
@@
In survey applications, LLMs can significantly improve testing pipelines by generating plausible test data Simmons and Hare (2023); Hämäläinen et al. (2023); Wang et al. (2023a). By simulating a variety of respondent behaviors and answers, LLMs allow the identification of weaknesses and biases in survey instruments. While not perfect, LLMs may provide diverse datasets that are helpful for testing and improving survey methods to ensure reliable data collection and analysis Namikoshi et al. (2024). However, in this case, too, it is important to note the potential mismatch between model-generated data and actual human responses Bisbee et al. (2023); von der Heyde et al. (2023); Hämäläinen et al. (2023).
6 Towards a Future of Evaluating AOV in LLMs
As we have discussed, evaluating AOV in LLMs offers opportunities alongside notable challenges (§5). To harness these opportunities while addressing the challenges, we show below key areas where focused action may lead to substantial improvements (WHAT to do?).
@@ 6 Towards a Future of Evaluating AOV in LLMs
The current methods for evaluating AOV in LLMs within the pipeline sometimes lack the necessary rigor for robust and reliable evaluations, especially due to the unstable results from the current evaluation methods. We call for the development of a more robust and fine-grained evaluation pipeline that can better capture the nuances of human-like expressions in LLM outputs. Besides, there is a great gap in the evaluation benchmarks. The current existing benchmarks for evaluating the opinions in LLMs such as OpinionQA Santurkar et al. (2023) and MMLU Hendrycks et al. (2021b) are static. Interactive benchmarks such as AlpacaEval Li et al. (2023c) and MT-Bench Zheng et al. (2023) focus more on general preferences. Therefore, more human-centered and fine-grained benchmarks from cognitive and social sciences should also be explored and extended to validate the “human” factors within the models in real-world scenarios.
@@
Understanding and improving the evaluation of AOV in LLMs requires insights from multiple disciplines. Interdisciplinary collaboration can provide a deeper understanding of both human cognitive processes and model behaviors. It is crucial to involve experts from different fields, e.g. survey methodology, psychology and sociology, to guide how we design and analyze the evaluations Dwivedi et al. (2023); Eckman et al. (2024). Research driven by interdisciplinary hypotheses can enhance our understanding of how well LLMs capture human-like AOV from a broader perspective.
@@ 7 Limitations
In this work, we present a survey and commentary on the progress and challenges of evaluating AOV in LLMs. There are several key limitations that should be acknowledged:
@@
This survey predominantly focuses on works with subjective context related to opinions and values. As a result, other relevant areas such as emotion detection, (e.g. Wang et al., 2023b; Li et al., 2023a), which might implicitly contain value expressions, have not been included here. Future research could explore a broader range of related works beyond AOV.
@@
The discussion on the evaluation pipeline in this work may be limited in scope, mainly focusing on the four evaluation stages, however decisions in each step have potential for profound impact on results. While we provide an overview of the evaluation pipeline with diverse approaches in each evaluation stage, there may be additional aspects or single features of the evaluation pipeline that were not thoroughly examined or highlighted, such as detailed pre-processing and data augmentation methods, intermediate representation analysis and error analysis methods. Future studies could delve deeper into these aspects to contribute in providing an even more comprehensive understanding of the evaluation process of AOV in LLMs.
@@
This work primarily focuses on the evaluation aspect of LLMs and does not extensively explore their detailed use cases in social science and society. While evaluating AOV in LLMs is undoubtedly important, it is equally crucial to consider how these models can be applied in various domains to address real-world challenges. Future research could explore the broader implications of LLMs in potential use cases, such as social science research, policy-making, education, and other societal applications to provide a more holistic perspective on their utility and impact.
@@ 8 Ethical Considerations
As discussed in §5.2, overpersonalizing and anthropomorphizing AI models might raise privacy risks and ethical concerns. The use of LLMs in social science research brings up important ethical questions regarding privacy, consent, and the potential for harm. Most LLMs are instruction-tuned with safety mechanisms to avoid sensitive topics and conflicts Grigis and De Angeli (2024). Despite this, researchers must exercise extreme caution due to the potential mismatch between LLM outputs and actual human opinions, which can also lead to harmful consequences due to misleading conclusions. To prevent these issues, it is crucial to continuously monitor and address cultural and value biases in LLM outputs, ensuring that AI usage does not perpetuate stereotypes or lead to unfair or harmful treatment of any group Aakanksha et al. (2024). Additionally, opinionated LLMs can influence users’ views and decision-making, necessitating careful monitoring and engineering Jakesch et al. (2023); Sharma et al. (2024). Researchers must remain vigilant and transparent about the limitations and ethical complexities of employing LLMs in their studies.
@@ Acknowledgments
We thank the members of SODA Lab and MaiNLP Lab from LMU Munich, and the members of the Social Data Science Group from University of Mannheim for their constructive feedback. Rob van der Goot from ITU Copenhagen provided valuable comments on an early draft. BM, ACH and FK are supported by Munich Center for Machine Learning (MCML). XW and BP are supported by ERC Consolidator Grant DIALECT 101043235 and in parts by Independent Research Fund Denmark (DFF) Sapere Aude grant 9063-00077B. TH is supported by Gates Cambridge Trust (grant OPP1144 from the Bill & Melinda Gates Foundation). MAH is supported by MCML.
@@ A.1 Overview of Surveyed Works
To compile this survey, we conducted a comprehensive review of recent literature on AOV in LLMs. We focused on identifying works that address these aspects, using keywords “attitude”, “opinion”, “value”, “culture”, “moral”, along with “LLMs”, “Language Models”. We utilized academic databases with a primary focus on *CL proceedings and Arxiv papers published from 2022 to the present (June 2024). Especially, we concentrated on evaluating and probing methods described in these papers.
We show the overview of a total of the 63 surveyed works in Table 1. The surveyed works are categorized into three main topics: Attitudes/Opinions, Values, and Others. The first two categories correspond to the main terms we defined in §2, each further subdivided into specific sub-topics. The additional category, Others, includes works that extend beyond the primary terms but still evaluate opinions and values in LLMs during their deployment. We categorize the topics into subtopics, as described in §3.3.

### v3

@@ 4.4 Evaluation
By drawing upon well-known measures of inter-annotator agreement and similarity measures, alignment metrics have been employed to measure the alignment of human and LLM responses. These measures include Cohen’s Kappa Argyle et al. (2023); Hwang et al. (2023), 1-Wasserstein distance (WD) Santurkar et al. (2023); Hwang et al. (2023); Sanders et al. (2023), Kullback–Leibler (KL) divergence Dominguez-Olmedo et al. (2023); Sun et al. (2024), the Euclidean distance Wang et al. (2023b), Jensen-Shannon Distance Durmus et al. (2024), Jaccard similarity Geng et al. (2024), as well as correlation and statistical analysis Kalinin (2023); Sun et al. (2024); Jiang et al. (2024). Moreover, metrics have been applied to measure the alignment between variables, such as regression models for measuring the correlations between single features of different personas Bisbee et al. (2023) and between different nations Benkler et al. (2023).
@@
Several studies conduct robustness evaluations on LLM outputs to validate the opinion consistency under different prompt perturbations or persona steerings, typically by measuring the percentage of samples that reach the same answer as a consistency score (e.g. Feng et al., 2023; Rao et al., 2023b; Röttger et al., 2024; Shu et al., 2024), or by measuring the entropy of answers (e.g. Dominguez-Olmedo et al., 2023; Tjuatja et al., 2024; Wang et al., 2024b). A few specific metrics for robustness evaluation have been proposed. Santurkar et al. (2023) introduce a consistency metric that measures whether LLMs align with the same group across various topics. Bonagiri et al. (2024) use Semantic Graph Entropy (SaGE) to evaluate LLM consistency, while Wright et al. (2024) employ average total variation distance (TVD) between model responses across different demographic categories to assess answer variability. Additionally, reliability and validity metrics in psychometric have been applied to measure the internal consistenty of the answers Shu et al. (2024).
5 Opportunities and Challenges in Evaluating AOVs in LLMs
Drawing from findings summarized in §3 and §4 from recent advances, we now focus on the methodological and practical perspectives regarding opportunities and challenges of evaluating AOVs in LLMs (WHERE?). This section addresses several key issues starting with the need to understand the models themselves (step 1), followed by the necessity for human-AI alignment (step 2), and finally, the implications from and for downstream applications in social sciences (step 3).
@@ 5.1 Understanding the Model
The essential discussion on the impact of evaluating AOVs in LLMs should start with the models themselves – the agents creating output. Our understanding of these models is limited (much like our understanding of ourselves as humans) Hassija et al. (2024). As studying how people respond to questions and express opinions helps us understand human behavior, examining how models do the same can enhance our knowledge of these models.
Evaluating AOVs Helps Understand Model Behavior.
By effectively evaluating AOVs in LLMs, we could potentially better explain their behavior in those subjective contexts, which could reveal why models produce certain opinions and values, helping us to better interpret their outputs. Apart from the textual output, tracking model internal behavior is also of interest, for example, to examine whether there exist skill neurons Wang et al. (2022); Voita et al. (2024). Investigating the internal working mechanisms of models enhances their interpretability, helping to make their operations more transparent and understandable. Currently, there is a lack of work linking AOVs evaluations to model interpretability. Addressing this gap would significantly contribute to the understanding and reliability of LLM outputs, especially in subjective contexts.
Evaluating AOVs Helps Understand Model Biases.
@@
One challenge in evaluating the output of LLMs is that the methods used can themselves be brittle. For example, in multiple-choice survey question settings, several studies rely on the first token logits (probabilities) of model output to map the options with the highest logits (as discussed in the Logits-Based Output section in §4.3). However, Wang et al. (2024a, b) observe that the first token logits do not always match the textual outputs and sometimes the mismatch rate can be over 50% in Llama2-7B Touvron et al. (2023) and Gemma-7B Team et al. (2024). A few works have also highlighted models being sensitive to option ordering Binz and Schulz (2023); Pezeshkpour and Hruschka (2024); Zheng et al. (2024); Shu et al. (2024); Wei et al. (2024), or generating inconsistent outputs to semantically equivalent situations Jang and Lukasiewicz (2023); Bonagiri et al. (2024); Wright et al. (2024), making evaluation unstable. Therefore, any evaluation for AOVs in LLMs should be accompanied by extensive robustness tests (see the Robustness Metrics section in §4.4). Wang et al. (2024a, b) propose to look at the text by training classifiers on the annotated LLM outputs, which typically requires a lot of human effort and may not be generalizable. Developing context-aware evaluation metrics to capture human-like nuances in LLM outputs is an ongoing research focus for model interpretability.
@@ 5.2 Human-AI Alignment
After understanding the model, aligning LLMs with human AOVs and ensuring that they perform safely and effectively is the next crucial phase.
@@
Alignment methods, such as Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022), focus on the problem of aligning LLMs to human values, which requires transferring the human values into alignment target for training and evaluating the models Klingefjord et al. (2024). However, current evaluations have often been coarse, highlighting the need for more fine-grained benchmarks to assess alignment effectively Lee et al. (2024b). One fundamental challenge RLHF faces is the problem of misspecification Casper et al. (2023). The diversity of human values cannot be easily represented by a single reward function. Current alignment evaluation benchmarks and reward model training rely on individual preference but lack consideration of the nature of diversity in human opinion. A more fine-grained evaluation of AOVs with respect to social choice Conitzer et al. (2024) or social awareness Yang et al. (2024) will help us better understand the alignment process and design a better socially-aware alignment algorithms Conitzer et al. (2024).
@@ 5.3 Implications from and for Social Science Applications
Considering the potential and challenges from the model perspective, we will now explore the feasibility of deploying AOVs in LLMs in downstream social science applications. LLMs, with their ability to process vast amounts of text data, could provide valuable insights into human values and behaviors. Again, caution must be exercised to address inherent biases and alignment issues that may arise.
@@
Currently, we have no means of aligning LLMs to accurately represent the diversity of human opinions necessary for reliable public opinion polling and similar tasks. The existing literature highlights numerous challenges, particularly in replicating non-US values Benkler et al. (2023); Arora et al. (2023); Simmons (2023); Rao et al. (2023a); Qu and Wang (2024). While some argue that LLM surveys might provide insights into hard-to-reach populations, the risk remains that these groups are difficult to model by LLMs von der Heyde et al. (2023); Namikoshi et al. (2024).
Human AOVs Help Evaluate AOVs in LLMs.
While there is a great gap between Human AOVs and those in LLMs, human-centered applications can enhance our understanding and validation of AOVs in LLMs. In survey methodology, responding to a survey question involves several cognitive steps, mainly including comprehension, retrieval, judgment, and reporting Tourangeau et al. (2000); Groves et al. (2004); Tourangeau (2018). Figure 2 illustrates a basic model of the human survey response process. Despite fundamental differences, the behavioral study of machines can benefit from that of animals Rahwan et al. (2019), as well as of humans Greasley and Owen (2016); van Dijk et al. (2023). Therefore, by integrating these human-centered cognitive processes into the examination of how LLMs respond to survey questions, we are able to gain valuable insights into the models and then modify the models to better align with human processes. Still, while concepts from human AOVs are certainly helpful in studying LLMs, we should also keep in mind at all times that they are after all not humans and should caution against the anthropomorphism we discussed in the previous section.
@@
In survey applications, LLMs can improve testing pipelines by generating plausible test data Simmons and Hare (2023); Hämäläinen et al. (2023); Wang et al. (2023a). By simulating a variety of respondent behaviors and answers, LLMs allow the identification of weaknesses and biases in survey instruments. While not perfect, LLMs may provide diverse datasets that are helpful for testing and improving survey methods to ensure reliable data collection and analysis Namikoshi et al. (2024). However, in this case, too, it is important to note the potential mismatch between model-generated data and actual human responses.
6 Towards a Future of Evaluating AOVs in LLMs
As we have discussed, evaluating AOVs in LLMs offers opportunities alongside notable challenges (§5). To harness these opportunities while addressing the challenges, we show below key areas where focused action may lead to substantial improvements (WHAT to do?).
@@ 6 Towards a Future of Evaluating AOVs in LLMs
The current methods for evaluating AOVs in LLMs within the pipeline sometimes lack the necessary rigor for robust and reliable evaluations, especially due to the unstable results from the current evaluation methods. We call for the development of a more robust and fine-grained evaluation pipeline that can better capture the nuances of human-like expressions in LLM outputs. Besides, there is a great gap in the evaluation benchmarks. The current existing benchmarks for evaluating the opinions in LLMs such as OpinionQA Santurkar et al. (2023) and MMLU Hendrycks et al. (2021) are static. Interactive benchmarks such as AlpacaEval Li et al. (2023d) and MT-Bench Zheng et al. (2023) focus more on general preferences. Therefore, more human-centered and fine-grained benchmarks from cognitive and social sciences should also be explored and extended to validate the “human” factors within the models in real-world scenarios.
@@
Understanding and improving the evaluation of AOVs in LLMs requires insights from multiple disciplines. Interdisciplinary collaboration can provide a deeper understanding of both human cognitive processes and model behaviors. It is crucial to involve experts from different fields, e.g. survey methodology, psychology and sociology, to guide how we design and analyze the evaluations Li et al. (2023c); Dwivedi et al. (2023); Eckman et al. (2024). Research driven by interdisciplinary hypotheses can enhance our understanding of how well LLMs capture human-like AOVs from a broader perspective.
@@ 7 Limitations
In this work, we present a survey and commentary on the progress and challenges of evaluating AOVs in LLMs. There are several key limitations that should be acknowledged:
@@
This survey predominantly focuses on works with subjective context related to opinions and values. As a result, other relevant areas such as emotion detection, (e.g. Wang et al., 2023b; Li et al., 2023a), which might implicitly contain value expressions, have not been included here. Future research could explore a broader range of related works beyond AOVs.
@@
The discussion on the evaluation pipeline in this work may be limited in scope, mainly focusing on the four evaluation stages, however decisions in each step have potential for profound impact on results. While we provide an overview of the evaluation pipeline with diverse approaches in each evaluation stage, there may be additional aspects or single features of the evaluation pipeline that were not thoroughly examined or highlighted, such as detailed pre-processing and data augmentation methods, intermediate representation analysis and error analysis methods. Future studies could delve deeper into these aspects to contribute in providing an even more comprehensive understanding of the evaluation process of AOVs in LLMs.
@@
This work primarily focuses on the evaluation aspect of LLMs and does not extensively explore their detailed use cases in social science and society. While evaluating AOVs in LLMs is undoubtedly important, it is equally crucial to consider how these models can be applied in various domains to address real-world challenges. Future research could explore the broader implications of LLMs in potential use cases, such as social science research, policy-making, education, and other societal applications to provide a more holistic perspective on their utility and impact.
@@ 8 Ethical Considerations
As discussed in §5.2, overpersonalizing AI models can raise privacy risks and ethical concerns. The use of LLMs in social science research can bring up important ethical questions regarding privacy, consent, and the potential for harm. While LLMs are instruction-tuned with safety mechanisms to avoid sensitive topics Grigis and De Angeli (2024), researchers must be cautious of mismatches between LLM outputs and human opinions, which can lead to misleading and harmful conclusions. To prevent these issues, it is crucial to continuously monitor and address cultural and value biases in LLM outputs, ensuring that AI usage does not perpetuate stereotypes or lead to unfair or harmful treatment of any group Aakanksha et al. (2024); Abdurahman et al. (2024). Additionally, opinionated LLMs can influence users’ views and decision-making, necessitating careful monitoring and engineering Jakesch et al. (2023); Sharma et al. (2024). Researchers must remain vigilant and transparent about the limitations and ethical complexities of employing LLMs in their studies.
@@ Acknowledgments
We thank the members of SODA Lab and MaiNLP labs from LMU Munich, and the members of the Social Data Science Group from University of Mannheim for their constructive feedback. Rob van der Goot from ITU Copenhagen provided valuable comments on an early draft. XW and BP are supported by ERC Consolidator Grant DIALECT 101043235. TH is supported by Gates Cambridge Trust (grant OPP1144 from the Bill & Melinda Gates Foundation).
@@ A.1 Overview of Surveyed Works
To compile this survey, we conducted a comprehensive review of recent literature on AOVs in LLMs. We focused on identifying works that address these aspects, using keywords “attitude”, “opinion”, “value”, “culture”, “moral”, along with “LLMs”, “Language Models”. We utilized academic databases with a primary focus on *CL proceedings and Arxiv papers published from 2022 to the present (primarily by June 2024, with partial updates in September 2024). In particular, we concentrated on the evaluation and probing methods described in these papers.
We show the overview of a total of the 67 surveyed works in Table 1. The surveyed works are categorized into three main topics: Attitudes/Opinions, Values, and Others. The first two categories correspond to the main terms we defined in §2, each further subdivided into specific sub-topics. The additional category, Others, includes works that extend beyond the primary terms but still evaluate opinions and values in LLMs during their deployment. We categorize the topics into subtopics, as described in §3.3.